# Extreme Conformal Prediction: Reliable Intervals for High-Impact Events

### v1

@@ 3.1 Unilateral prediction intervals
instead of Eq. 5, where $\hat{Q}_{1-\alpha}(\bm{x})$ is the pretrained quantile regression model at level $1-\alpha$. Let $y_{\text{min}}\in\mathbb{R}\cup\{-\infty\}$ be the lower endpoint of the distribution of $Y$ (or of the conditional distribution $Y\mid\bm{X}=\bm{x}$, if known). Then, following the usual procedure, the resulting interval is
$C(\bm{x})=\left(y_{\text{min}},\hat{Q}_{1-\alpha}(\bm{x})+\hat{q}_{\alpha}\right],$ (8)
where $\hat{q}_{\alpha}$ is, in classical conformal prediction, the $(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c})$-quantile of the calibration scores $s_{1},\ldots,s_{n_{c}}$, i.e., the order statistic $s_{(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil)}$.
@@ 3.2 Calibrative extrapolation
$\mathbb{P}(S>y)=\mathbb{P}(S>u)\mathbb{P}(S>y\mid S>u)\approx\mathbb{P}(S>u)\left(1+\xi\frac{y-u}{\sigma(u)}\right)_{+}^{-1/\xi},\quad\forall y\geq u,$ (10)
where $\xi\in\mathbb{R}$ and $\sigma(u)>0$ are the shape and scale parameters and $u$ is an intermediate threshold. Under very mild assumptions on the distribution $F_{S}$ of $S$, this approximation is theoretically justified as $u$ tends to the upper endpoint of $F_{S}$  (Balkema and de Haan, 1974; Pickands III, 1975). In practice, $u$ is typically chosen as the empirical $\tau_{0}=(1-k/n_{c})$-quantile $\hat{Q}_{\tau_{0}}$ of $S$ for some $k<n_{c}$. The tuning parameter $k$ is the number of exceedances used for estimation of the parameters $\sigma$ and $\xi$, for instance, by maximum likelihood. Quantiles of $S$ can then be estimated at probability levels beyond the data range using the approximation
$\hat{Q}^{\text{GPD}}_{\tilde{\tau}}:=\hat{Q}_{\tau_{0}}+\frac{\hat{\sigma}}{\hat{\xi}}\left[\left(\dfrac{1-\tau_{0}}{1-\tilde{\tau}}\right)^{\hat{\xi}}-1\right],\quad\forall\tilde{\tau}>\tau_{0}.$ (11)
Theoretical guarantees of these estimators typically require that $k\to\infty$ and $k/n_{c}\to 0$ as $n_{c}\to\infty$ to ensure the correct trade-off between bias and variance; see de Haan and Ferreira (2006) for more details.
Asymptotically as $n_{c}\to\infty$ and under additional second-order conditions, using $\hat{q}^{e}_{\alpha}:=\hat{Q}^{\text{GPD}}_{1-\alpha}$ would satisfy Eq. 9 with equality (in a suitable limiting sense). However, on a finite calibration sample and with a fixed threshold $u$, $\hat{Q}^{\text{GPD}}_{1-\alpha}$ can underestimate the true quantile due to estimation and approximation biases, respectively (Roodman, 2018; Bücher and Zhou, 2021; Zeder et al., 2023). We, therefore, follow here a more conservative approach. Alternatively to choosing $\hat{q}^{e}_{\alpha}$ as $\hat{Q}^{\text{GPD}}_{1-\alpha}$, we may use the upper endpoint of a $(1-\alpha_{2})$-confidence interval for ${F}^{-1}_{S}(1-\alpha_{1})$, the $(1-\alpha_{1})$-quantile of the calibration scores $S$, for two suitable levels $\alpha_{1},\alpha_{2}\in(0,1)$. The following proposition shows that, if the upper endpoint of the confidence interval has correct coverage, then the resulting extreme conformal prediction interval satisfies Eq. 1.
@@
Let $\alpha_{1},\alpha_{2}\in(0,1)$, and $[Q_{L},Q_{U}]$ be a $(1-\alpha_{2})$-confidence interval for ${F}^{-1}_{S}(1-\alpha_{1})$, the $(1-\alpha_{1})$-quantile of the calibration scores $S$. If the confidence interval has correct coverage, i.e. $\mathbb{P}(U_{L}\leq{F}^{-1}_{S}(1-\alpha_{1})\leq U_{Q})\geq 1-\alpha_{2}$, and if
@@
Let $[Q_{L},Q_{U}]$ be the $(1-\alpha_{2})$-confidence interval for $q:={F}^{-1}_{S}(1-\alpha_{1})$. Then, by assumption,
@@
Two natural choices for $\alpha_{1}$ and $\alpha_{2}$ that satisfy Eq. 12 are $\alpha_{1}=\alpha_{2}=\alpha/2$, which is analogous to a Bonferroni correction (Bonferroni, 1936), and $\alpha_{1}=\alpha_{2}=1-(1-\alpha)^{1/2}$, which is analogous to the Šidák correction (Šidák, 1967). Although the Šidák correction is slightly less conservative, the difference is negligible for small values of $\alpha$.
@@
There are several well-studied approaches for obtaining extreme quantile confidence intervals (CI) using the GPD approximation (Coles, 2001; Davison and Hinkley, 1997; Davison et al., 2003; de Haan and Zhou, 2022), including the profile likelihood CI, bootstrap CI and delta method CI with normal approximation. The profile likelihood CI typically represents uncertainty best and yields the most conservative CI upper endpoint. However, with small sample sizes when $\alpha$ is very small, estimating the upper endpoint of the CI can be challenging since the derivative of the right side of the profile log-likelihood is close to zero. This sometimes makes finding the crosspoint between the profile curve and the chi-square confidence line numerically difficult. This may lead to numerically undefined or infinite CI upper endpoints. The bootstrap approach comes in different variations, both for the sampling step (nonparametric, parametric) and for the aggregation step (basic, percentile, normal, etc…). We here focus on the nonparametric bootstrap with percentile aggregation, as there are the most commonly used. The bootstrap can give reliable confidence intervals but might be less conservative than the profile-likelihood approach. Finally, the delta method CI is computationally less expensive than the other two alternatives, but it provides intervals that are asymmetric around the quantile estimate, which is not realistic for large quantiles. Moreover, it can also suffer from numerical instability issues, due to its matrix inversion step, and fail to yield meaningful CIs.
@@ 4.1 Experimental setup
$\begin{cases}\bm{X}\sim\mathcal{U}\left([-1,1]^{10}\right),\\ Y\mid\bm{X}=\bm{x}\sim\sigma(\bm{x})\cdot\varepsilon_{Y},\end{cases}$ (13)
@@
with $\sigma(\bm{x}):=1+6\phi(x_{1},x_{2})$, where $\phi$ is the bivariate Gaussian density with correlation $0.9$. We consider two scenarios for the noise variable $\varepsilon_{Y}$: a heavy-tailed Student $t$ distribution $t_{\alpha(\bm{x})}$, with covariate-dependent tail index $\alpha(\bm{x})=1/\xi(\bm{x}):=7\cdot\{1+\exp(4x_{1}+1.2)\}^{-1}+3$, and a light-tailed Gaussian $\mathcal{N}(0,1)$ distribution. The former choice corresponds exactly to the generating process used in two extreme quantile regression benchmark studies (Velthoen et al., 2023; Pasche and Engelke, 2024).
@@
We consider several sizes for the calibration sets, with $n_{c}\in\{10^{3},10^{3.5},10^{4}\}$ observations. For each calibration size, we repeat the experiments 100 times. We consider extreme PI confidence levels $1-\alpha$, with $\alpha\in\{10^{-3},10^{-3.5},10^{-4},10^{-4.5},10^{-5}\}$. We consider two choices for the initial “pretrained” quantile predictions: the conditional-quantile ground truth and a pretrained extreme quantile regression model. The former aims at assessing the methods with ideal initial predictions. As all conformalization methods are translation invariant, adding first-order bias to the pretrained model would always lead to the same final PIs. For the latter choice, we use the extreme quantile regression neural networks (EQRN) model, as it performed best on this benchmark dataset (Pasche and Engelke, 2024). It is pretrained on 5,000 observations generated from Eq. 13. Its hyperparameters and architecture were selected based on validation GPD deviance with a grid search. For each calibration size, repetition, confidence level, and initial predictions, we perform the following conformalization procedures and assess their average population coverage on a separate test set of $10^{6}$ observations.
  • GPD profile: $\hat{q}^{e}_{\alpha}$ is the endpoint of the profile-likelihood CI for the calibration-score quantile.
  • GPD bootstrap: $\hat{q}^{e}_{\alpha}$ is the endpoint of the nonparametric bootstrap percentile CI for the calibration-score quantile.
  • GPD delta: $\hat{q}^{e}_{\alpha}$ is the endpoint of the delta method CI for the calibration-score quantile.
@@
  • GPD simple: $\hat{q}^{e}_{\alpha}$ is the GPD $(1-\alpha)$-quantile estimate of the calibration scores.
  • Classical: Classical (split, single-sided) conformalized quantile regression based on the calibration-score order statistics.
@@ 4.2 Coverage results
Figure 1 shows the distribution of the computed test coverage for each considered conformalization method, confidence level, and calibration size, for the Student $t$ noise and ground truth original predictions. The chosen confidence levels are particularly large relative to the size of the calibration sets. Hence, for most scenarios, $\alpha<1/(n_{c}+1)$. In those cases, the classical conformalization method always yields trivial infinite intervals. They have a coverage of $1$, but are uninformative and of no practical use. On the other hand, the other methods, relying on peaks over threshold extrapolation instead of empirical quantiles, are able to yield finite PIs, even when $\alpha\ll 1/(n_{c}+1)$.
@@
The simple GPD estimates of the calibration-score $(1-\alpha)$-quantile do not seem to provide sufficient coverage with small calibration samples and for the larger confidence levels, likely due to the GPD estimation error or approximation biases. The other three methods, relying on the confidence intervals for the score quantiles (and Proposition 3.1), achieve much better coverage and, in most cases, satisfy Eq. 1 as their coverage is larger than $1-\alpha$ on average. However, those GPD CI-based methods seem consistently overconservative for lower confidence levels.
In general, the profile likelihood method seems the most conservative, compared to the nonparametric bootstrap and delta method alternatives, as anticipated. It also satisfies the coverage guarantee in all scenarios. Its downside is the numerical difficulty, described in Section 3.2. With the implementation at hand, this issue arose, in the worst case, in $85\%$ of the repetitions for $n_{c}=1000$ and the lowest $\alpha$ value, but quickly decreased to, at most, $2\%$ for $n_{c}=3163$ and $0\%$ for $n_{c}=10000$. This instability is understandable in the former truly extreme case, as the confidence level is more than two orders of magnitude larger than the level for which PIs are obtainable with the classical conformal method and as the likelihood only relies on $50$ observations to estimate a $(1-5\cdot 10^{-6})$-confidence interval for a $(1-5\cdot 10^{-6})$-quantile.
@@
Figure A.1 in Appendix A.1 shows the same coverage distribution when the data-generating process has light-tailed Gaussian noise, instead of heavy-tailed Student $t_{4}$ noise. In comparison, all methods tend to result in significantly more conservative intervals in terms of the coverage. In particular, all three GPD CI-based approaches always result in more coverage than necessary.
Figure 2 shows the coverage distributions for the EQRN predictions and Student $t_{4}$ noise. We observe that, although being accurate predictions of the conditional quantile, in terms of integrated squared error (Pasche and Engelke, 2024), the EQRN predictions, in this case, undercover when considered as a PI endpoint. The conformalization results closely match those for the ground-truth quantile predictions, although the coverage seems smaller for the largest confidence levels, for all methods. The scenario with $\alpha=10^{-5}$ and the largest sample size is the only one for which the GPD profile approach seems to slightly undercover. All the other non-infinite alternatives also undercover for this largest confidence level. The extreme conformal prediction methods all outperform the original EQRN prediction in all scenarios, except for the GPD simple on small calibration sizes.
@@
As a takeaway, our practical recommendation for conservative and informative high-confidence PIs is to use the profile-likelihood method if its estimate is finite and the bootstrap-based estimate otherwise. This combination results in the profile-likelihood conservativeness, in most scenarios, and avoids the potential infinite intervals with the bootstrap estimate, which is still conservative enough in the majority of scenarios, in cases of numerical difficulties. We call this method the “GPD safeprofile” PI. Alternatively, considering the maximum of the bootstrap and delta-method PI endpoints as a replacement for unstable profile likelihood situations could be more conservative but might be less stable.

### v2

@@ 3.1 Single-sided prediction intervals
instead of Eq.˜5, where $\hat{Q}_{1-\alpha}(\bm{x})$ is the pretrained quantile regression model at level $1-\alpha$. Let $y_{\text{min}}\in\mathbb{R}\cup\{-\infty\}$ be the lower endpoint of the distribution of $Y$ (or of the conditional distribution $Y\mid\bm{X}=\bm{x}$, if known). Then, following the usual procedure, the resulting interval is
$\hat{C}(\bm{x})=\left(y_{\text{min}},\;\hat{Q}_{1-\alpha}(\bm{x})+\hat{q}_{\alpha}\right],$ (8)
where $\hat{q}_{\alpha}$ is, in classical conformal prediction, the $\{\left\lceil(n_{c}+1)(1-\alpha)\right\rceil/n_{c}\}$-quantile of the calibration scores $S_{1},\ldots,S_{n_{c}}$, i.e., the order statistic $S_{(\left\lceil(n_{c}+1)(1-\alpha)\right\rceil)}$.
@@ 3.2 Calibrative extrapolation
$\mathbb{P}(S>y)=\mathbb{P}(S>u)\mathbb{P}(S>y\mid S>u)\approx\mathbb{P}(S>u)\left\{1+\xi\frac{y-u}{\sigma(u)}\right\}_{+}^{-1/\xi},\quad y\geq u,$ (10)
where $\xi\in\mathbb{R}$ and $\sigma(u)>0$ are the shape and scale parameters and $u$ is an intermediate threshold. Under very mild assumptions on the distribution $F_{S}$ of $S$, this approximation is theoretically justified as $u$ tends to the upper endpoint of $F_{S}$ (GPDBalkema; GPDPickands). In practice, $u$ is typically chosen as the empirical $\tau_{0}=(1-k/n_{c})$-quantile $\hat{Q}_{\tau_{0}}$ of $S$ for some $k<n_{c}$. The tuning parameter $k$ is the number of exceedances used for estimation of the parameters $\sigma$ and $\xi$, for instance, by maximum likelihood. Quantiles of $S$ can then be estimated at probability levels beyond the data range using the approximation
[…]
Theoretical guarantees of these estimators typically require that $k\to\infty$ and $k/n_{c}\to 0$ as $n_{c}\to\infty$ to ensure the correct trade-off between bias and variance; see DF2006 for more details.
Asymptotically, as $n_{c}\to\infty$ and under additional second-order conditions, using $\hat{q}^{e}_{\alpha}:=\hat{Q}^{\text{GPD}}_{1-\alpha}$ would satisfy Eq.˜9 with equality (in a suitable limiting sense as $\tau_{0}\to 1$). But, since the calibration sample is finite and the level $\tau_{0}$ fixed, $\hat{Q}^{\text{GPD}}_{1-\alpha}$ can underestimate the true quantile due to estimation and approximation biases, respectively (Roodman2018; BMvsPOT; ZederPasche23). We, therefore, follow here a more conservative approach. Alternatively to choosing $\hat{q}^{e}_{\alpha}$ as $\hat{Q}^{\text{GPD}}_{1-\alpha}$, we may use the upper endpoint of a $(1-\alpha_{2})$-confidence interval for ${F}^{-1}_{S}(1-\alpha_{1})$, the $(1-\alpha_{1})$-quantile of the calibration scores $S$, for two suitable levels $\alpha_{1},\alpha_{2}\in(0,1)$. The following proposition shows that, if the upper endpoint of the confidence interval has correct coverage, then the resulting extreme conformal prediction interval satisfies Eq.˜1.
@@
Let $\alpha_{1},\alpha_{2}\in(0,1)$, and $Q_{L},Q_{U}]$ be a $(1-\alpha_{2})$-confidence interval for ${F}^{-1}_{S}(1-\alpha_{1})$, the $(1-\alpha_{1})$-quantile of the calibration scores $S$. If the confidence interval has correct coverage, i.e. $\mathbb{P}\{U_{L}\leq{F}^{-1}_{S}(1-\alpha_{1})\leq U_{Q}\}\geq 1-\alpha_{2}$, and if
@@
Let $Q_{L},Q_{U}]$ be the $(1-\alpha_{2})$-confidence interval for $q:={F}^{-1}_{S}(1-\alpha_{1})$. Then, by assumption,
@@
A natural choice for $\alpha_{1}$ and $\alpha_{2}$ that satisfies Eq.˜12 with equality is $\alpha_{1}=\alpha_{2}=1-(1-\alpha)^{1/2}$, which is analogous to the Šidák correction (Sidak1967). Another notable choice is $\alpha_{1}=\alpha_{2}=\alpha/2$, which is analogous to a Bonferroni correction (bonferroni1936). Although the latter is, in this case, overconservative, the difference is negligible for small values of $\alpha$.
@@
There are several well-studied approaches for obtaining extreme quantile confidence intervals (CI) using the GPD approximation (IExtr; DavisonBootstrap; DavisonBootstrap2; HaanZhouExtrBoot), including the profile likelihood, the bootstrap and the normal-approximation delta method. The profile-likelihood CI typically represents the inherently asymmetrical uncertainty best and yields the most conservative CI upper endpoint. For very small $\alpha$ and small sample sizes, it sometimes suffers from numerical difficulties for estimating the CI’s upper endpoint. They arise as the derivative of the profile log-likelihood can get close to zero on its right-hand side, which sometimes makes finding the crosspoint between the profile curve and the chi-square confidence line numerically difficult. Slightly varying the value of the GPD threshold can sometimes solve the issue. The bootstrap approach comes in different variations, both for the sampling step (nonparametric, parametric) and for the aggregation step (basic, percentile, normal, etc.). We here consider the nonparametric bootstrap with percentile aggregation, as it is the most commonly used. The bootstrap can give reliable confidence intervals, but, compared to the profile-likelihood approach, its upper endpoint estimates might not always be conservative enough. Finally, the delta-method CI is computationally less expensive than the other two alternatives, but it provides intervals that are symmetric around the quantile estimate, which is not realistic for large quantiles. Moreover, it can also suffer from numerical instability issues, due to its matrix inversion step, and fail to yield meaningful CIs. These alternative extreme CI methods are further discussed and compared in Section˜4.
To summarise, when given a high confidence level $1-\alpha$, above or close to $n_{c}/(n_{c}+1)$, our proposed extreme conformal prediction interval is
$\hat{C}^{e}(\bm{x})=\left(y_{\text{min}},\;\hat{Q}^{e}_{1-\alpha}(\bm{x})+\hat{q}^{e}_{\alpha}\right],$ (13)
where $\hat{Q}^{e}_{1-\alpha}(\cdot)$ is a prefitted extreme quantile regression model and $\hat{q}^{e}_{\alpha}$ is the upper endpoint of an appropriate GPD profile-likelihood confidence interval for a high quantile of the calibration nonconformity scores, as discussed above. If a two-sided extreme PI is preferred, one can use the classical nonconformity scores from Eq.˜5, instead of Eq.˜7, to obtain
$\hat{C}^{e}(\bm{x})=\left[\hat{Q}^{e}_{\alpha/2}(\bm{x})-\hat{q}^{e}_{\alpha},\;\hat{Q}^{e}_{1-\alpha/2}(\bm{x})+\hat{q}^{e}_{\alpha}\right],$ (14)
where a second prefitted extreme quantile regression model $\hat{Q}^{e}_{\alpha/2}(\cdot)$ is required for extrapolating the lower-tail conditional quantiles.
3.3 Extensions to other conformal approaches and nonexchangeable data
We introduce our extreme conformal prediction approach as a variant of the well-established split conformalized quantile regression, as its use of base quantile predictions naturally results in varying-length intervals, which, in addition to valid marginal coverage, can also achieve good conditional coverage (ConformalizedQR; ConformalIntrRev), and as the split procedure is significantly less computationally costly than full-conformal or $k$-fold alternatives. Nevertheless, it is, in principle, adaptable to other conformal approaches, including different base regression models and nonconformity scores, the alternative full-conformal and $k$-fold procedures, and weighted methods for nonexchangeable data.
The extension to different base predictive models and scores is the most straightforward, as the extreme conformalization described in Section˜3.2 is agnostic to their definitions. This includes the classical split-conformal approach, using a conditional-mean base regression model $\hat{\mu}(\bm{x})$, instead of the quantile predictions $\hat{Q}_{\alpha/2}(\bm{x}),\hat{Q}_{1-\alpha/2}(\bm{x})$ (SplitConformalMargProof; papadopoulos2008inductive; Lei18ConformalPred), and its heteroscedastic variants (papadopoulos2008normalizedCoform; papadopoulos2011regressionknn; Lei18ConformalPred). Although these alternatives would not require extreme quantile regression, they either yield fixed-length PIs or tend to underestimate conditional variability (ConformalizedQR). The extension to full-conformal procedures (vovk1999ml; vovk2005algorithmic; shafervovk2008tutorial) is also straightforward, but requires repeating both the predictive model fit and our extreme conformalization from Section˜3.2 for a dense grid of potential response values. Although it makes a more efficient use of the data than the split variant, it has an extremely high computational cost. The same applies, but less extremely, to $k$-fold approaches, such as Jackknife+/CV+ and cross-conformal prediction (barber2021predictive; vovk2015cross). The extensions of our method to these alternative conformalisation approaches are described in more detail in Appendix˜A.1.
Weighted conformal approaches allow for relaxing the exchangeability assumption to account, for example, for distribution shifts or drifts. In those approaches, each calibration observation is assigned a weight $w_{i}\in[0,1]$, $i=1,\ldots,n_{c}$, which reflects its similarity to the test observation, or more generally, the similarity between its score $S_{i}$ and the test score $S_{\text{test}}$. The classical empirical quantile of the calibration scores $\hat{q}_{\alpha}$ is then replaced by a weighted sample quantile. More precisely, it is redefined as the $(1-\alpha)$-quantile of the weighted empirical distribution $\bar{w}^{n}\sum_{i=1}^{n_{c}}w_{i}\delta_{s_{i}}+\bar{w}^{n}\delta_{\infty}$, where $\bar{w}^{n}=(\sum_{i=1}^{n_{c}}w_{i}+1)^{-1}$ and $\delta_{x}$ is the Dirac measure at $x$ (Barber2023nonstatConform; TibshiraniCoformShift). As our proposed $\hat{q}^{e}_{\alpha}$ from Section˜3.2 is, for extreme confidence levels, based on likelihood inference, the natural analogous extension to nonexchangeable data would be to rely on weighted likelihood inference instead. Using the sample weights $w_{i}\in[0,1]$, $i=1,\ldots,n_{c}$, the procedure would remain the same as in Section˜3.2, only using the weighted
$\ell^{\text{GPD}}_{w}(\sigma,\xi)=-\log(\sigma)\sum_{i=1}^{n_{c}}w_{i}-\left(\frac{1}{\xi}+1\right)\sum_{i=1}^{n_{c}}w_{i}\log\left(1+\xi\frac{s_{i}-u}{\sigma}\right)_{+},$ (15)
instead of the classical GPD log-likelihood, to infer the $(1-\alpha_{2})$-CI endpoint for ${F}^{-1}_{S_{\text{test}}}(1-\alpha_{1})$. This weighted alternative should achieve a similar effect as classical weighted conformal prediction methods: proportionally use the calibration scores most similar to the test point during conformalization, to correct for nonexchangeable distribution drifts and shifts. This weighted extension is further discussed and applied in Section˜5.
@@ 4.1 Experimental setup
$\begin{cases}\bm{X}\sim\text{Unif}\left([-1,1]^{10}\right),\\ Y\mid\bm{X}=\bm{x}\sim\sigma(\bm{x})\cdot\varepsilon_{Y},\end{cases}$ (16)
@@
with $\sigma(\bm{x}):=1+6\phi(x_{1},x_{2})$, where $\phi$ is the bivariate Gaussian density with correlation $0.9$. We consider two scenarios for the noise variable $\varepsilon_{Y}$: a heavy-tailed Student $t$ distribution $t_{\alpha(\bm{x})}$, with covariate-dependent tail index $\alpha(\bm{x})=1/\xi(\bm{x}):=7\cdot\{1+\exp(4x_{1}+1.2)\}^{-1}+3$, and a light-tailed Gaussian $\text{N}(0,1)$ distribution. The former choice corresponds exactly to the generating process used in two extreme quantile regression benchmark studies (gbex; Pasche2024).
@@
We consider several sizes for the calibration sets, with $n_{c}\in\{10^{3},10^{3.5},10^{4}\}$ observations. For each calibration size, we repeat the experiments 100 times. We consider extreme PI confidence levels $1-\alpha$, with $\alpha\in\{10^{-3},10^{-3.5},10^{-4},10^{-4.5},10^{-5}\}$. We consider three choices for the base quantile predictions $\hat{Q}^{e}_{1-\alpha}(\cdot)$: the conditional-quantile ground truth and two different pretrained extreme quantile regression models. The ground-truth choice aims at assessing the methods with ideal initial predictions. As all conformalization methods are translation invariant, adding first-order bias to the pretrained model would always lead to the same final PIs. For the first pretrained extreme quantile predictions, we use the extreme quantile regression neural networks (EQRN) model, as it performed best on this benchmark dataset (Pasche2024), aiming at assessing our procedure with accurate but realistic initial predictions. The EQRN model is pretrained on 5,000 observations generated from Eq.˜16. Its hyperparameters and architecture were selected based on validation GPD deviance with a grid search. Lastly, to investigate the performance of our extreme conformalization procedure for a poorly-performing model, which could happen in practice due to poor historical modelling choices, we also consider a linear GPD quantile model as base predictions. More precisely, we use linear quantile regression to obtain a conditional threshold $\hat{u}(\bm{x})$, and model the exceedances with a GPD having a linear parametrization $\sigma=\sigma_{0}+\sigma_{1}x_{1}+\sigma_{2}x_{2}$ for its scale parameter. Although having the GPD extrapolation potential, its linear parametrization is a bad fit for the highly nonlinear dependence of $Y$ on $\bm{X}$.
For each calibration size, repetition, confidence level, and initial predictions, we perform the following conformalization procedures and assess their average population coverage on a separate test set of $10^{6}$ observations. The methods compared for conformalization are:
  • GPD profile: $\hat{C}^{e}(\bm{x})$ from Eq.˜13, using the GPD profile-likelihood CI for $\hat{q}^{e}_{\alpha}$.
  • GPD bootstrap: $\hat{C}^{e}(\bm{x})$, using the GPD nonparametric bootstrap percentile CI for $\hat{q}^{e}_{\alpha}$.
  • GPD delta: $\hat{C}^{e}(\bm{x})$, using the GPD delta-method CI for $\hat{q}^{e}_{\alpha}$.
@@
  • GPD simple: $\hat{C}^{e}(\bm{x})$, but using the simple GPD $(1-\alpha)$-quantile estimate of the calibration scores instead of the endpoint of a CI for $\hat{q}^{e}_{\alpha}$.
  • Classical: Single-sided version of the classical split conformalized quantile regression (still using the extreme quantile predictions $\hat{Q}^{e}_{1-\alpha}(\cdot)$, for a fair comparison).
@@ 4.2 Coverage results
Figure˜1 shows the distribution of the computed test coverage for each considered conformalization method, confidence level, and calibration size, for the Student $t$ noise and ground truth original predictions. The chosen confidence levels are particularly large relative to the size of the calibration sets. Hence, for most scenarios, $\alpha<1/(n_{c}+1)$. In those cases, the classical conformalization method always yields trivial infinite intervals. They have a coverage of $1$, but are uninformative and of no practical use. On the other hand, the other methods, relying on peaks over threshold extrapolation instead of empirical quantiles, are able to yield finite PIs, even when $\alpha\ll 1/(n_{c}+1)$.
@@
The simple GPD estimates of the calibration-score $(1-\alpha)$-quantile do not seem to provide sufficient coverage with small calibration samples and for the larger confidence levels, likely due to the GPD estimation error or approximation biases. The other three methods, relying on the confidence intervals for the score quantiles (and Proposition˜3.1), achieve much better coverage and, in most cases, satisfy Eq.˜1 as their coverage is larger than $1-\alpha$ on average. However, those GPD CI-based methods seem consistently overconservative for lower confidence levels.
In general, the profile likelihood method seems the most conservative, compared to the nonparametric bootstrap and delta method alternatives, as anticipated. It also satisfies the coverage guarantee in all scenarios. Its downside is the numerical difficulty, described in Section˜3.2. With the implementation at hand, this issue arose, in the worst case, in $85\%$ of the repetitions for $n_{c}=1000$ and the lowest $\alpha$ value, but quickly decreased to, at most, $2\%$ for $n_{c}=3163$ and $0\%$ for $n_{c}=10000$. This instability is understandable in the former truly extreme case, as the confidence level is more than two orders of magnitude larger than the level for which PIs are obtainable with the classical conformal method and as the likelihood only relies on $50$ observations to estimate a $(1-5\cdot 10^{-6})$-confidence interval for a $(1-5\cdot 10^{-6})$-quantile. However, this instability issue appears to often be resolved by slightly varying the choice of the GPD threshold, or of the $\alpha_{1}$ and $\alpha_{2}$ split.
@@
Figure˜A.1 in Appendix˜A.2 shows the same coverage distribution when the data-generating process has light-tailed Gaussian noise, instead of heavy-tailed Student $t_{4}$ noise. In comparison, all methods tend to result in significantly more conservative intervals in terms of the coverage. In particular, all three GPD CI-based approaches always result in more coverage than necessary.
Figure˜2 shows the coverage distributions for the EQRN predictions and Student $t_{4}$ noise. We observe that, although being accurate predictions of the conditional quantile, in terms of integrated squared error (Pasche2024), the EQRN predictions, here, undercover when considered as a PI endpoint. Note that even very accurate quantile predictions are still likely to lead to undercoverage, as a local quantile underestimation typically leads to a larger coverage loss than the coverage gain from a local overestimation of the same amplitude, due to the generally decreasing probability density in the tails. The conformalization results closely match those for the ground-truth quantile predictions, although the coverage seems smaller for the largest confidence levels, for all methods. The scenario with $\alpha=10^{-5}$ and the largest sample size is the only one for which the GPD profile approach seems to slightly undercover. All the other finite alternatives also undercover for this largest confidence level. The CI-based extreme conformal prediction methods all outperform the original EQRN prediction in all scenarios.
@@
Figure˜A.2 in Appendix˜A.2 shows the same coverage distributions when using the poorly-fitting linear GPD quantile predictions, described in Section˜4.1, instead of the EQRN model. The base predictions severely undercover, when used as PIs. We observe that, even with poor quantile modelling choices, in all considered scenarios, our extreme conformalization approach results in a mean test coverage always greater than the desired levels, in this case for all three GPD CI-based variants. Interestingly, the intervals seem overall more marginally conservative than with the EQRN base predictions. Thus, even with this choice of a poorly-fit base model, undoubtingly worsening local adaptivity and coverage, our method still results in nontrivial PIs consistently achieving sufficient marginal coverage.
As a takeaway, our practical recommendation for conservative and informative high-confidence PIs is to use the profile-likelihood version of our method. In case it suffers from numerical issues and varying the GPD threshold or the $\alpha_{1}$ and $\alpha_{2}$ split slightly does not solve them (or if those are desired fixed), the bootstrap-based CI can be used instead. This combination results in the profile-likelihood conservativeness, in most scenarios, and avoids the cases of potential numerical difficulties by using the bootstrap estimate, which is still conservative enough in the majority of scenarios. We call this method the “GPD safeprofile” PI. Alternatively, considering the maximum of the bootstrap and delta-method PI endpoints as a replacement in unstable profile likelihood situations could be more conservative but might be less stable.